OpenStudy (anonymous):

solve for m in v = √F/mk - u^2 *square root is over all of the equation after the equals sign

OpenStudy (anonymous):

\(v=\sqrt{\frac{F}{mk}-u^2}\) like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

square both sides \[v^2=\frac{F}{mk}-u^2\]

OpenStudy (anonymous):

add \(u^2\) to both sides. flip both sides multiply both sides by \(\frac{F}{k}\)

OpenStudy (anonymous):

thank you!